# LLM-based Resource-Oriented Intention Inference for Static Resource Detection

### v2

@@ 3.3. Static Resource Leak Detection
3.3.1. Enumerating Control-Flow Paths
Similar to most existing resource leak detection methods (Torlak and Chandra, 2010; Wu et al., 2016b; fin, 2023; inf, 2023), our detection approach is also built upon control-flow analysis. We construct control-flow graphs (CFGs) for the code to detect and extract pruned control-flow paths from these graphs.
Given a method-level code snippet, we begin by parsing it into an abstract syntax tree (AST) and then proceed to construct a CFG by traversing the AST nodes and inserting various control-flow edges between the code statements. The procedure is implemented based on the top of an established CFG construction tool (pro, 2023). Within the CFG, certain nodes, such as if-statements and switch-statements, give rise to multiple subsequent control-flow branches. These particular nodes are referred to as branch-nodes within the CFG. Next, we enumerate the paths within the CFG, considering the paths between the entry-node and exit-node. To ensure the efficiency of the enumeration process, we employ two pruning operations.
Loop Structures. Loop statements in the CFG can introduce circular paths, leading to infinite-length paths during the enumeration. To prevent this, we adopt a strategy that involves expanding the true-branch of a loop only once. For instance, for a while-loop, we visit the statement in its true-branch once and then proceed directly to the end of the loop.
Resource-Independent Branches. The presence of branches in the CFG can exponentially increase the number of paths during enumeration, but most of the branches might not be related to resource management. To address this, for a branch-node, we further examine whether all of its branches are independent of resource acquisition/release and procedural exit (e.g, return-statements). If the conditions are met, we only retain one branch for the branch-node, as all the other branches exhibit the same resource acquisition/release behaviors. The examination involves the inferred resource-oriented intentions, which help determine whether a CFG node acquires/releases resources.
@@ 3.3.1. Enumerating Control-Flow Paths
3.3.2. Detecting Resource Leaks
@@
For each concerned resource involved in the inferred intentions, the detection process involves a two-stage path analysis, as illustrated in Algorithm 1. This algorithm takes the concerned resource $res$, all the control-flow paths $\mathcal{P}$, and the resource-oriented intention set $\mathcal{I}$ as inputs. The first stage (lines 1-16) is a single-path analysis that initially identifies leak-risky paths of the concerned resource based on the $ACQUIRE$ and $RELEASE$ intentions. The second stage (lines 17-29) is a cross-path analysis that greedily eliminates the false-alarm-introducing branch-nodes by considering the $VALIDATE$ intentions. After completing both stages, the algorithm (lines 31-36) checks if any path has a corresponding status of $true$. If such a path exists, a resource leak of $res$ is reported.
@@
9:      else if $RELEASE(res,ln)\in\mathcal{I}$ then
@@
Stage 1: Single-Path Analysis. The algorithm conducts an iterative process, sequentially processing each $path$ in $\mathcal{P}$ and traversing the nodes within $path$ (lines 2-17). During this node traversal, a descriptor counter $rd\_counter$ keeps track of the $ACQUIRE$ and $RELEASE$ intentions involving the resource $res$ within $path$ (lines 3-11). If $rd\_counter$ is greater than 0, it indicates that $path$ is a potential leak-risky path for $res$, and as a result, the risky status of $path$ (i.e., $path.risky$) is initialized as $true$ (lines 12-16).
@@
Stage 2: Cross-Path Analysis. As outlined in Section 2, the presence of branches, denoted as unreachable branches, where the resource is inaccessible, gives rise to the possibility of false-positive (FP) leak-risky paths. This, in turn, results in false alarms being reported at line 32. During this phase, we conduct cross-path analysis utilizing the inferred reachability validation intentions to eliminate the false alarms.
@@
Intuitively, addressing this problem involves specifically analyzing each branch-node intended for reachability validation and identifying its unreachable branch. However, the task is complicated by the diverse implementations of validation APIs, as discussed in Section 2. Some APIs (e.g., isClosed()) return true to indicate that the resources are reachable, while others (e.g., isActive()) indicate the opposite. To overcome this challenge, our approach utilizes a novel heuristic that collectively analyzes all branches of a branch-node to assess whether it introduces false alarms. If affirmative, such branch-nodes are designated as false-alarm-introducing (FAI) nodes. These FAI nodes are characterized by its one branch being completely leak-risky, while the others are completely not leak-risky. A branch is considered completely leak-risky when all the paths belonging to it exhibit a risky status of $true$. The completely leak-risky branch is the unreachable branch, and the paths within it are FP leak-risky paths for $res$. The insight behind this decision is that if the other branch (the completely not leak-risky branch) is the unreachable one, it suggests the presence of more severe functional bugs rather than resource leaks, such as null pointer errors.
@@
To implement this heuristic, our algorithm initiates by identifying all branch-nodes in the CFG and iterates through them from back to front (lines 18-30). During this process, each branch-node $br\_node$ is examined to determine if it involves a $VALIDATE$ intention. If not, the algorithm skips $br\_node$, recognizing that it does not influence the risky status of the paths (lines 21-24). On the contrary, when $br\_node$ involves a $VALIDATE$ intention, the algorithm conducts two grouping operations (line 24 and line 26, respectively). This results in the division of paths containing $br\_node$ into branches $(B_{1},B_{2},...)$. All these branches share a common path prefix (a subpath from the entry-node to $br\_node$) but belong to different branches of $br\_node$. This implies that, at a specific execution path to $br\_node$, $(B_{1},B_{2},...)$ represent potential subsequent branches of $br\_node$. Upon the branches, the algorithm performs the heuristic by executing the Propagate procedure and updating the risky status of FP leak-risky paths to $false$ (line 28).
@@
4. Implementation
The current implementation of InferROI utilizes the widely adopted gpt-4 (the version is gpt-4-0613) as the LLM and sets the model temperature as 0 to remove generation randomness. During the static resource leak detection, we take into account certain Java-specific language features, such as try-(catch)-finally structures and try-with-resources statements. Regarding the try-(catch)-finally structure, the statements within the finally block are executed in all control-flow paths. As a result, when enumerating the control-flow paths on control-flow graphs (CFGs), we pay special attention to processing the statements within the finally block to ensure the paths are complete. For the try-with-resources statement, the acquired resources are automatically closed, and there is no need for manual resource closure (try, 2023). To address this, we incorporate a simple rule-based post-processing step to filter out any false alarms introduced by the try-with-resources statements.
5. Evaluation
We conduct a comprehensive set of experiments to evaluate the effectiveness of InferROI in both inferring resource-oriented intentions and detecting resource leaks. To accomplish this, we gather 86 bugs from the DroidLeaks dataset (Liu et al., 2019), resulting in a collection of 172 (2 $\times$ 86) code snippets, each comprising both the buggy version and the corresponding fixed version. InferROI typically takes an average of 3 to 8 seconds to process each code snippet.
In the initial phase of evaluation (RQ1), we manually annotate the resources and resource-oriented intentions present in the 172 code snippets and subsequently measure the extent to which InferROI successfully infers these intentions. Next, we apply InferROI to detect resource leaks in the 172 code snippets, comparing its detection rate and false alarm rate against those of baseline detectors (RQ2). By doing so, we assess whether InferROI can surpass the performance of the baseline detectors in accurately identifying resource leaks. In the third part of our evaluation (RQ3), we explore the efficacy of InferROI in detecting resource leaks when applied to real-world open-source projects. This investigation enables us to gauge InferROI’s performance under real-world conditions. Finally, we conduct an ablation study (RQ4) to analyze and compare the contributions of two key components of InferROI.
@@ 5. Evaluation
  • RQ1 (Effectiveness in Intention Inference): To what degree can InferROI accurately infer the resource-oriented intentions in code?
@@
  • RQ2 (Effectiveness in Leak Detection): How effectively can InferROI detect resource leaks in the DroidLeaks dataset? Can it outperform baseline detectors in this regard?
@@
  • RQ3 (Open-Source Project Scanning): How effective is InferROI in detecting previously-unknown resource leak bugs when applied to open-source projects?
@@
  • RQ4 (Ablation Study): What is the contribution of the intention inference and static analysis in InferROI?
5.1. Intention Inference (RQ1)
In this evaluation, we utilize the DroidLeaks dataset (Liu et al., 2019) to evaluate the effectiveness of InferROI’s LLM-based resource-oriented intention inference.
5.1.1. Data
The DroidLeaks dataset consists of resource leak bugs from open-source Android applications by analyzing commits dedicated to fixing such bugs. Each bug in the dataset is associated with a concerned resource type, which is determined by reviewing the corresponding bug-fixing commit. For instance, if a commit addresses a resource leak involving Cursor, the concerned resource type for that bug is Cursor. By retrieving the bug-fixing commit and the previous commit, both the buggy and fixed versions are obtained for each bug.
In the original DroidLeaks dataset, 116 bugs were considered for evaluating the performance of eight widely-used resource leak detectors, namely Code Inspection (cod, 2023b), Infer (inf, 2023), Lint (lin, 2023), FindBugs (fin, 2023), Relda2-FS (Wu et al., 2016a), RElda2-FI (Wu et al., 2016a), Elite (Liu et al., 2016), and Verifier (Vekris et al., 2012). For our evaluation, we attempt to obtain all the 116 bugs and successfully obtain both the buggy and fixed versions for 86 of them (some commits are unavailable due to application maintenance issues). Thus, we conduct experiments based on the 86 bugs (experimented bugs), resulting in a collection of 172 code snippets encompassing 28 different concerned resource types, such as Cursor and InputStream.
5.1.2. Procedure
To evaluate the effectiveness of InferROI’s intention inference, two authors independently annotate all three types of resource-oriented intentions present in the 172 code snippets. In cases of disagreement, they engage in discussions to reach a final decision, ensuring consistent and accurate annotations. These manual annotations serve as the ground truth ($Ground$) for the evaluation.
Subsequently, we apply InferROI to the code snippets to infer the intentions (denoted as $Inferred$), and then calculate the precision and recall of the inference based on the manual annotations. Precision represents the proportion of intentions inferred by InferROI that align with the ground truth, while recall indicates the proportion of true intentions from the ground truth that InferROI correctly infers. The precision and recall calculations are as follows:
$\text{precision}=\frac{|Ground\cap Inferred|}{|Inferred|};\text{recall}=\frac{|Ground\cap Inferred|}{|Ground|}$
Furthermore, we evaluate the resource coverage of InferROI and compare it with the eight detectors (i.e., baselines) evaluated in the DroidLeaks dataset. For InferROI, we determine the portion of the 19 concerned resource types covered by the inferred intentions; for the eight baseline detectors, we calculate the resource coverage directly using the data from the original DroidLeaks dataset.
5.1.3. Results
InferROI achieves a precision of 74.6% and a recall of 81.8% in resource-oriented intention inference, indicating its ability to accurately and comprehensively identify intentions within the code snippets. Table 1 presents the resource coverage of InferROI and the eight compared baseline detectors. Notably, InferROI demonstrates a higher ability to cover more concerned resource types involved in the DroidLeaks dataset. We have observed that certain resource types, such as Camera and MediaPlayer, remain uncovered in our intention inference process. These resources are typically managed by the Android lifecycle and are not explicitly present in the code of the buggy and fixed versions.
Through analysis of the evaluation results and the experimentation with various prompts, we attribute the effectiveness of InferROI’s intention inference to two main factors. Firstly, the powerful capabilities of LLMs in leveraging background knowledge and understanding code context allow them to exhibit strong advantages in inferring resource-oriented intentions. For example, InferROI can identify the AndroidHttpClient-related intentions, which are ignored by other baseline detectors. Secondly, our well-designed prompt template plays a critical role in appropriately eliciting the relevant capabilities of LLMs. When we experimented with a prompt template lacking the first two leading instructions (see Section 3.2.1), the inference precision and recall experienced a decline of 20-30%. This underscores the importance of prompt engineering for LLMs, e.g., guiding them with clear instruction-chains to enhance their reasoning abilities in intention inference.
Summary: InferROI demonstrates a remarkable precision of 74.6% and a commendable recall of 81.8% in resource-oriented intention inference. This achievement can be attributed to the potent capabilities of LLMs, coupled with the efficacy of our designed prompt template.
Table 1. Resource Coverage Across the DroidLeaks Dataset
Detector # Covered Resource Types (Resource Coverage)
Code Inspection 16 (57.1%)
Infer 11 (39.3%)
Lint 3 (10.7%)
FindBugs 9 (32.1%)
Relda2-FS 6 (21.4%)
RElda2-FI 6 (21.4%)
Elite 2 (7.1%)
Verifier 2 (7.1%)
InferROI (ours) 19 (67.9%)
5.2. Resource Leak Detection (RQ2)
@@
In this evaluation, we assess the effectiveness of InferROI in detecting resource leaks using the DroidLeaks dataset.
5.2.1. Procedure
@@
We use the 86 bugs collected from the DroidLeaks dataset (as described in Section 5.1.1) as the benchmark. Following DroidLeaks, we adopt the metrics bug detection rate and false alarm rate to evaluate the detection effectiveness of InferROI.
For each bug in our experiments, we apply InferROI to its buggy version and fixed version. If InferROI reports a resource leak in the buggy version related to the corresponding concerned resource type, it is considered to have correctly detected the bug. Conversely, if InferROI reports a resource leak in the fixed version for the corresponding concerned resource type, it signifies a false alarm. The bug detection rate and false alarm rate are calculated as follows:
bug detection rate=# detected bugs in buggy versions# total experimented $\textit{bug detection rate}=\frac{\textit{\# detected bugs in buggy versions}}{\textit{\# total experimented bugs}}$
false alarm rate=# reported false alarms in fixed versions# total experimented $\textit{false alarm rate}=\frac{\textit{\# reported false alarms in fixed versions}}{\textit{\# total experimented bugs}}$
To compare the performance of InferROI with the eight detectors (i.e., baselines) evaluated in DroidLeaks, we calculate the bug detection rate and false alarm rate for each baseline directly using the data from the original DroidLeaks dataset. For fair comparison, we use the same set of the 86 bugs as the experimented bugs for both InferROI and all the baseline detectors, instead of specifically selecting different experimented bugs for each detector as done in the original DroidLeaks evaluation.
5.2.2. Results
The evaluation results are presented in Table 2. Notably, InferROI achieves a high bug detection rate of 53.5% and a low false alarm rate of 8.1% on the 86 bugs.
Among the baseline detectors, only Code Inspection has a higher bug detection rate than InferROI, with a relative difference of 6.5% (3 more detected bugs). However, Code Inspection’s false alarm rate is significantly higher than InferROI, with a relative difference of 414.3% (29 more false alarms). On the other hand, five detectors (Lint, FindBugs, RElda2-FI, Elite, and Verifier) demonstrate lower false alarm rates than InferROI, with relative differences ranging from 42.9% to 100% (3 to 6 fewer false alarms). However, their bug detection rates are significantly lower than InferROI, with relative differences ranging from 76.1% to 93.5% (35 to 43 fewer detected bugs). In comparison, Infer achieves the best balance between the two evaluation metrics (i.e., detection rate versus false alarms) among all the baseline detectors. Yet, InferROI outperforms Infer with a 24.3% relatively higher bug detection rate (9 more detected bugs) and a 56.3% relatively lower false alarm rate (9 fewer false alarms).
An important aspect worth highlighting is that InferROI demonstrates promising detection performance based solely on single methods, without relying on complex interprocedural program analysis, such as call graphs and aliasing. This can be attributed to the advantages of the resource-oriented intention inference utilized by InferROI, as discussed in the insights provided in Section 2. However, it is essential to acknowledge that the resource-oriented intention inference cannot entirely replace the value of more robust program analysis techniques. For instance, in the case of buggy versions not detected by InferROI, a substantial portion is caused by the intricacies of Android’s lifecycle management and callback mechanisms. This observation indicates an important future direction, i.e., to explore better integration of LLM-based resource-intention inference with more advanced program analysis techniques, which could enable more general and powerful resource leak detection.
Summary: InferROI achieves a notable bug detection rate of 53.5% and a commendably low false alarm rate of 8.1% on the DroidLeaks dataset. This remarkable performance shows a promising potential of detecting resource leaks based on resource-oriented intentions rather than purely relying on mechanical matching of predefined resource acquisition/release APIs and null-checking conditions.
Table 2. Resource Leak Detection on the DroidLeaks Dataset
@@ 5.2.2. Results
Lint 10 (11.6%) 0 (0.0%)
FindBugs 6 (6.9%) 0 (0%)
Relda2-FS 11 (12.7%) 9 (10.4%)
RElda2-FI 8 (9.3%) 4 (4.6%)
Elite 6 (6.9%) 4 (4.6%)
Verifier 3 (3.4%) 2 (2.3%)
InferROI (ours) 46 (53.5%) 7 (8.1%)
5.3. Open-Source Project Scanning (RQ3)

### v3

@@ III-C Static Resource Leak Detection
III-C1 Enumerating Control-Flow Paths
Similar to most existing resource leak detection methods [3, 4, 6, 7], our detection approach is also built upon control-flow analysis. We construct control-flow graphs (CFGs) for the code to detect and extract pruned control-flow paths from these graphs.
Given a method-level code snippet, we begin by parsing it into an abstract syntax tree (AST) and then proceed to construct a CFG by traversing the AST nodes and inserting various control-flow edges between the code statements. The procedure is implemented based on the top of an established CFG construction tool [24]. Within the CFG, certain nodes, such as if-statements and switch-statements, give rise to multiple subsequent control-flow branches. These particular nodes are referred to as branch-nodes within the CFG. Next, we enumerate the paths within the CFG, considering the paths between the entry-node and exit-node. To ensure the efficiency of the enumeration process, we employ two pruning operations.
Loop Structures. Loop statements in the CFG can introduce circular paths, leading to infinite-length paths during the enumeration. To prevent this, we adopt a strategy that involves expanding the true-branch of a loop only once.
Resource-Independent Branches. The presence of branches in the CFG can exponentially increase the number of paths during enumeration, but most of the branches are not related to resource management. To address this, for a branch-node, we further prune its branches that are independent of resource acquisition/release and procedural exit (e.g, return-statements).
@@ III-C1 Enumerating Control-Flow Paths
III-C2 Detecting Resource Leaks
@@
For each concerned resource involved in the inferred intentions, the detection process involves a two-stage path analysis, as illustrated in Algorithm 1. This algorithm takes the concerned resource $res$, all the control-flow paths $\mathcal{P}$, and the resource-oriented intention set $\mathcal{I}$ as inputs. The first stage (lines 1-16) is a single-path analysis that initially identifies leak-risky paths of the concerned resource based on the $ACQ$ and $REL$ intentions. The second stage (lines 17-29) is a cross-path analysis that greedily eliminates the false-alarm-introducing branch-nodes by considering the $VAL$ intentions. After completing both stages, the algorithm (lines 31-36) checks if any path has a corresponding status of $true$. If such a path exists, a resource leak of $res$ is reported.
@@
9:      else if $REL(res,ln)\in\mathcal{I}$ then
@@
Stage 1: Single-Path Analysis. The algorithm conducts an iterative process, sequentially processing each $path$ in $\mathcal{P}$ and traversing the nodes within $path$ (lines 2-17). During this node traversal, a descriptor counter $rd\_counter$ keeps track of the $ACQ$ and $REL$ intentions involving the resource $res$ within $path$ (lines 3-11). If $rd\_counter$ is greater than 0, it indicates that $path$ is a potential leak-risky path for $res$, and as a result, the risky status of $path$ (i.e., $path.risky$) is initialized as $true$ (lines 12-16).
@@
Stage 2: Cross-Path Analysis. As outlined in Section II, the presence of branches, denoted as unreachable branches, where the resource is inaccessible, gives rise to the possibility of false-positive (FP) leak-risky paths. This, in turn, results in false alarms being reported at line 32. During this phase, we conduct cross-path analysis utilizing the inferred reachability validation intentions to eliminate the false alarms.
@@
Intuitively, addressing this problem involves specifically analyzing each branch-node intended for reachability validation and identifying its unreachable branch. However, the task is complicated by the diverse implementations of validation APIs, as discussed in Section II. Some APIs (e.g., isClosed()) return true to indicate that the resources are reachable, while others (e.g., isActive()) indicate the opposite. To overcome this challenge, our approach utilizes a novel heuristic that collectively analyzes all branches of a branch-node to assess whether it introduces false alarms. If affirmative, such branch-nodes are designated as false-alarm-introducing (FAI) nodes. These FAI nodes are characterized by its one branch being completely leak-risky, while the others are completely not leak-risky. A branch is considered completely leak-risky when all the paths belonging to it exhibit a risky status of $true$. The completely leak-risky branch is the unreachable branch, and the paths within it are FP leak-risky paths for $res$. The insight behind this decision is that if the other branch (the completely not leak-risky branch) is the unreachable one, it suggests the presence of more severe functional bugs rather than resource leaks, such as null pointer errors.
@@
To implement this heuristic, our algorithm initiates by identifying all branch-nodes in the CFG and iterates through them from back to front (lines 18-30). During this process, each branch-node $br\_node$ is examined to determine if it involves a $VAL$ intention. If not, the algorithm skips $br\_node$, recognizing that it does not influence the risky status of the paths (lines 21-24). On the contrary, when $br\_node$ involves a $VAL$ intention, the algorithm conducts two grouping operations (line 24 and line 26, respectively). This results in the division of paths containing $br\_node$ into branches $(B_{1},B_{2},...)$. All these branches share a common path prefix (a subpath from the entry-node to $br\_node$) but belong to different branches of $br\_node$. This implies that, at a specific execution path to $br\_node$, $(B_{1},B_{2},...)$ represent potential subsequent branches of $br\_node$. Upon the branches, the algorithm performs the heuristic by executing the Propagate procedure and updating the risky status of FP leak-risky paths to $false$ (line 28).
@@
IV Implementation
The current implementation of InferROI utilizes the widely adopted GPT-4 as the LLM and sets the model temperature as 0 to remove generation randomness. During the static resource leak detection, we take into account certain Java-specific language features, such as try-with-resources statements. For the try-with-resources statement, the acquired resources are automatically closed, and there is no need for manual resource closure [25]. To address this, we incorporate a simple rule-based post-processing step to filter out any false alarms introduced by the try-with-resources statements.
V Evaluation
We conduct a comprehensive set of experiments to evaluate the effectiveness of InferROI in detecting resource leaks by employing the DroidLeaks dataset [8] and the JLeaks dataset [13]. InferROI typically takes an average of 3 to 8 seconds to process each code snippet.
First, we apply InferROI to detect resource leaks in the buggy-version and fixed-version code snippets in the datasets, comparing its bug detection rate and false alarm rate against those of baseline detectors (RQ1). By doing so, we assess whether InferROI can surpass the performance of the baseline detectors in accurately identifying resource leaks. In the next part of our evaluation (RQ2), we explore the efficacy of InferROI in detecting resource leaks when applied to real-world open-source projects. This investigation enables us to gauge InferROI’s performance under real-world conditions. Finally, we manually annotate the resource-oriented intentions presented in code to measure the extent to which InferROI successfully infers these intentions (RQ3.a), and conduct an ablation study to analyze and compare the contributions of two key components of InferROI (RQ3.b).
@@ V Evaluation
  • RQ1 (Effectiveness in Leak Detection): How effectively can InferROI detect resource leaks in the DroidLeaks and JLeaks datasets? Can it outperform baseline detectors in this regard?
@@
  • RQ2 (Effectiveness in Project Scanning): How effective is InferROI in detecting previously-unknown resource leak bugs when applied to open-source projects?
  • RQ3.a (Accuracy in Intention Inference): To what degree can InferROI accurately infer the resource-oriented intentions in code?
@@
  • RQ3.b (Ablation Study): What is the contribution of the intention inference and static analysis in InferROI?
V-A Effectiveness in Resource Leak Detection (RQ1)
@@
In this evaluation, we assess the effectiveness of InferROI in detecting resource leaks using two datasets.
V-A1 Data
We utilize two datasets, namely DroidLeaks [8] and JLeaks [13], constructed by analyzing commits dedicated to fixing resource leak bugs. Each bug in the datasets is associated with a concerned resource type, which is determined by reviewing the corresponding bug-fixing commit. By retrieving the bug-fixing commit and the previous commit, both the buggy and fixed versions are obtained for each bug.
  • DroidLeaks: This dataset comprises resource leak bugs sourced from open-source Android applications. Originally, the DroidLeaks dataset included 116 bugs used to assess the performance of eight widely-used resource leak detectors [15, 7, 16, 6, 17, 18, 19]. In our evaluation, we attempt to acquire all 116 bugs, successfully obtaining both the buggy and fixed versions for 86 of them. Consequently, the final collection comprises 172 code snippets covering 28 different resource types, such as Cursor.
  • JLeaks: This is a more recent dataset containing resource leak bugs from open-source Java projects. The original JLeaks dataset included 1,094 bugs used to evaluate the performance of three widely-used resource leak detectors [6, 7, 26]. For our evaluation, we filtered out instances based on two conditions: 1) the resource type is annotated, and 2) both the buggy and fixed versions are free of syntax errors. This filtering process resulted in 784 bugs (1,568 code snippets) encompassing 368 different resource types.
V-A2 Baselines
We compare InferROI with three well-maintained industrial static detectors, as evaluated in previous studies [8, 13].
  • SpotBugs [6]: SpotBugs, formerly known as FindBugs, is a static analysis tool that identifies bugs and potential issues in Java programs.
  • Infer [7]: Infer is a static analysis tool developed by Facebook that detects various types of programming errors in C, C++, and Java code.
  • PMD [26]: PMD is a widely-used static code analysis tool that identifies potential issues in code.
V-A3 Procedure
@@
Following previous study [8], we adopt the metrics Bug Detection Rate and False Alarm Rate to evaluate the detection effectiveness of InferROI and baseline detectors.
TABLE I: Resource Leak Detection on DroidLeaks Dataset
@@
PMD 10 (11.6%) 6 (7.0%)
InferROI (ours) 51 (59.3%) 16 (18.6%)
TABLE II: Resource Leak Detection on JLeaks Dataset
Detector # Detected Bugs (Bug Detection Rate)
SpotBugs* 44 (24.3%)
Infer* 5 (9.4%)
PMD 341 (43.5%)
InferROI (ours) 508 (64.8%)
    Infer and SpotBugs encounter compilation issues with the majority of test instances.
For each bug in our experiments, we apply InferROI to its buggy version and fixed version. If InferROI reports a resource leak in the buggy version related to the corresponding concerned resource type, it is considered to have correctly detected the bug. Conversely, if InferROI reports a resource leak in the fixed version for the corresponding concerned resource type, it signifies a false alarm. The two metrics are calculated as follows:
Bug Detection Rate=# detected bugs in buggy versions# total experimented $\textit{Bug Detection Rate}=\frac{\text{\# detected bugs in buggy versions}}{\text{\# total experimented bugs}}$
False Alarm Rate=# reported false alarms in fixed versions# total experimented $\textit{False Alarm Rate}=\frac{\text{\# reported false alarms in fixed versions}}{\text{\# total experimented bugs}}$
V-A4 Results
The evaluation results for DroidLeaks and JLeaks are presented in Table I and Table II, respectively.
DroidLeaks: InferROI demonstrates a bug detection rate of 59.3% with a false alarm rate of 18.6% across 86 bugs in DroidLeaks. Compared to the baseline detectors, InferROI achieves notably higher bug detection, detecting 14~45 more bugs, while maintaining a comparable false alarm rate.
JLeaks: InferROI showcases a bug detection rate of 64.8% with a false alarm rate of 24.0% across 784 bugs in JLeaks. In contrast to the baseline detectors, InferROI achieves significantly superior bug detection, identifying 167~503 more bugs in JLeaks. It’s important to note that we do not report the false alarm rate of the baselines due to the unavailability of data necessary for its calculation in current JLeaks dataset.
Analysis: (i) Bug Detection Analysis: The higher bug detection rate of InferROI can be attributed to the advantages of the resource-oriented intention inference utilized by InferROI, as discussed in the insights provided in Section II (see RQ3.a for the resource coverage assessment). However, it is essential to acknowledge that the resource-oriented intention inference cannot entirely replace the value of more sound program analysis techniques. For instance, in the case of buggy versions not detected by InferROI, a substantial portion is caused by the intricacies of Android’s lifecycle management and callback mechanisms. This observation indicates an important future direction, i.e., to explore better integration of LLM-based resource-intention inference with more advanced program analysis techniques, which could enable more general and powerful resource leak detection. (ii) False Alarm Analysis: False alarms reported by InferROI primarily stem from incorrect or missing intentions. When InferROI incorrectly identifies an acquisition intention or overlooks a release intention, a false alarm occurs. To address this issue, we intend to explore the integration of LLM-based prompt engineering techniques, such as in-context learning and chain-of-thought [27], to enhance the capabilities of LLMs. Notably, a significant portion of these “false alarms” arises from misclassifications due to the quality of the ground truth. Since both datasets are curated based on bug-fixing commits, certain “fixed” versions still contain resource leaks. This underscores the ongoing challenge of ensuring the quality of resource leak datasets.
Summary: InferROI demonstrates a bug detection rate of 59.3% with a false alarm rate of 18.6% across 86 bugs in DroidLeaks, and a bug detection rate of 64.8% with a false alarm rate of 24.0% across 784 bugs in JLeaks. This performance shows a promising potential of detecting resource leaks based on resource-oriented intentions.
V-B Effectiveness in Open-Source Project Scanning (RQ2)